Pottery Blossom Inc. has been manufacturing its own finials for its curtain rods. The company is currently operating at 100% of capacity, and variable manufacturing overhead is charged to production at the rate of 63% of direct labor cost. The direct materials and direct labor cost per unit to make a pair of finials are $4 and $5, respectively. Normal production is 32,100 curtain rods per year.
A supplier offers to make a pair of finials at a price of $13.35 per unit. If Pottery Blossom accepts the supplier’s offer, all variable manufacturing costs will be eliminated, but the $49,400 of fixed manufacturing overhead currently being charged to the finials will have to be absorbed by other products.
(a) Prepare the incremental analysis for the decision to make or buy the finials. (Enter negative amounts using either a negative sign preceding the number e.g. -45 or parentheses e.g. (45).)

Make Buy Net Income
Increase (Decrease)

Direct materials

$______________ $______________ $______________

Direct labor

$______________ $______________ $______________

Variable overhead costs

$______________ $______________ $______________

Fixed manufacturing costs

$______________ $______________ $______________

Purchase price

$______________ $______________ $______________

Total annual cost

$______________ $______________ $______________

(b) Should Pottery Blossom buy the finials?

(Yes or No) ______, Pottery Blossom should ( buy or not buy ) __________ the finials.

(c) Would your answer be different in (b) if the productive capacity released by not making the finials could be used to produce income of $44,920?

(Yes or No) _______, income would select between _________ (increase or decrease) decreaseincrease by $enter net income in dollars
Answer #1

a)

Make Buy Net income Increase (Decrease)
Direct materials (32100*$4)= $128400 - $128400
Direct labor (32100*$5)= 160500 - 160500
Variable overhead costs (32100*$5*63%)= 101115 - 101115
Fixed overhead costs 49400 49400 0
Purchase price - (32100*$13.35)= $428535 -428535
Total annual cost $439415 $477935 $-38520

b) No, Pottery Blossom should not buy the finials as the total annual cost of buying is more than making the curtain rods.

c)

Make Buy Net income Increase (Decrease)
Direct materials (32100*$4)= $128400 - $128400
Direct labor (32100*$5)= 160500 - 160500
Variable overhead costs (32100*$5*63%)= 101115 - 101115
Fixed overhead costs 49400 49400 0
Opportunity cost 44920 - 44920
Purchase price - (32100*$13.35)= $428535 -428535
Total annual cost $484335 $477935 $6400

Yes, income would increase by $6400.